OpenStudy (anonymous):

Please help!Solve each system. In your work, identify the shape of the graph of each equation. 7y^2 +x^2=64 x+y=4

OpenStudy (shamil98):

x = -y + 4 substitute for x and solve for y. it's a circle and a line btw.

OpenStudy (anonymous):

so 7y^2 +(-y +4)^2=64?

OpenStudy (shamil98):

\[\huge 7y^2 + (4-y)^2 = 64\] \[\huge 7y^2 + 16 - 8y + y^2 = 64\] \[\huge 8y^2 - 8y - 48 = 0\]

OpenStudy (anonymous):

wait what does huge mean lol?

OpenStudy (shamil98):

I wrote this in latex, it just enlarges the font, if you're using another browser/phone/tablet it might not show...

OpenStudy (shamil98):

8(y^2 - y - 6) = 0 (y-3)(y+2) = 0 y = 3 y = -2 x + 3 = 4 x = 1 x - 2 = 4 x = 6 (1,3) and (6,-2) are your solutions

OpenStudy (anonymous):

Oh ok thank you!

OpenStudy (anonymous):

Wait is it a circle or an ellipse? And do I need to say there is a line passing through it?

OpenStudy (shamil98):

It's a circle with a line passing through it

OpenStudy (anonymous):

Could you help on one more? if not, that is okay